# The Generalization Ridge: Information Flow in Natural Language Generation

### v3

@@ 5.1 Predictive Information: Information Peaks at Intermediate Layers
Figure 2: Evolution of predictive information $I(Z_{\ell};Y)$, with lighter curves indicating later epochs. Each curve exhibits a three-phase trend: early layers rise, mid layers peak, and late layers decline.
@@
Information Peak (intermediate layers). $I(Z_{\ell};Y)$ continues to rise through the mid-to-upper layer, typically peaking before the final few blocks. For the GPT-2 Small on Synthetic Arithmetic dataset, the peak reaches $I\!\approx\!0.33$ in layers 10-11, coinciding with a jump from $\approx 0\%$ to $73\%$ ID accuracy and $57\%$ OOD accuracy. These layers appear to play a critical role in synthesizing abstract features that are essential for generalization.
@@
Layer $\mathbf{I(Z;Y)}$ Test Accuracy (%) All In-Dist. Out-Dist. Layer 1 0.0782 0.00 0.00 0.00 Layer 2 0.0776 0.00 0.00 0.00 Layer 3 0.0907 0.00 0.00 0.00 Layer 4 0.1156 0.00 0.00 0.00 Layer 5 0.1222 0.00 0.00 0.00 Layer 6 0.1275 0.00 0.00 0.00 Layer 7 0.1552 0.75 0.00 1.50 Layer 8 0.2594 19.45 0.00 40.70 Layer 9 0.2937 20.60 0.60 42.00 Layer 10 0.3295 64.25 72.90 56.90 Layer 11 0.3296 71.65 100.00 42.90 Layer 12 0.1501 70.45 100.00 41.40
@@
(a) 8 layers: Monotonic Increase
@@
Figure 3: Truncating GPT–2 to 8 layers removes the MI peak; 9–layer variants begin to exhibit a shallow peak, and the full 12–layer model shows a pronounced decline.
We empirically vary the depth of GPT-2 Small models fine-tuned on the CLUTRR dataset to examine how model capacity shapes information dynamics, shown in Figure 3. When the model is truncated to 8 layers, $I(Z;Y)$ increases monotonically—indicating insufficient capacity to develop an information peak. Upon increasing the depth to 9 layers, a shallow peak emerges, signifying the threshold at which the model begins to distinguish generalized features from memorized signals. The full 12-layer model exhibits a clear peak followed by a decline, confirming that the generalization ridge emerges only beyond a certain capacity threshold. These results underscore the role of architectural depth in information dynamics and sufficient capacity is required for the generalization ridge to emerge.
@@
This discussion aims to provide an intuitive illustration of our generalization ridge hypothesis, highlighting how attention to task-relevant signal tokens shifts across layers (the generalizable information). Specifically, we computed average signal attention across layers, identified the layer with peak signal attention, and compared it to final-layer signal and last-token attention, results are in Table 2. Signal tokens are defined as task-relevant tokens that the model must focus on to solve the task—for the synthetic dataset, these are tokens that appear after the character ‘S’; for CLUTRR, kinship-related entities; and for ECQA, the token corresponding to the correct answer option. This allows us to quantify where in the network attention to semantically important information is concentrated. Our findings show a pattern that supports our hypothesis: (1) Mid-to-late layers peak in signal attention, coinciding with the predictive information ridge (Figure 2), indicating where generalizable representations are strongest. (2) Final layers show reduced attention to signal tokens, suggesting a shift toward memorization to specific data point rather than predictive information abstraction. For example, in Qwen-2.5-0.5B on ECQA, signal attention peaks at Layer 17 (0.2458) but drops to 0.0066 in the final layer, where last-token attention dominates (0.1697).
@@
GPT-2 Small Synthetic 0.0227 0.0483 10 (0.0758) 0.0573 0.0400
GPT-2 Medium CLUTRR 0.0086 0.0155 19 (0.0453) 0.0147 0.0889
Qwen-2.5-0.5B ECQA 0.0225 0.0257 17 (0.2458) 0.0066 0.1697
LLaMA-3.1-8B ECQA 0.0220 0.0379 21 (0.1276) 0.0168 0.3904
@@ 5.2 Incremental Information Gain: Information Concentrates at Intermediate Layers
$z^{(\ell)}=z^{(\ell-1)}+\beta_{\ell}\cdot\text{block}^{(\ell)}(z^{(\ell-1)}),\qquad\beta_{\ell}\in\mathbb{R}_{\geq 0}.$
@@ E. Drive
Each synthetic sample consists of a sequence of 10 symbolic elements, where each element is formatted as S{signal}_N{noise}. The signal values follow an arithmetic progression modulo $K$, and the noise values are independently sampled from a uniform distribution with a fixed range of 100. During training, the modulus $K$ is set to 13. For evaluation, test sequences are generated using values of $K$ from the range $[5,26]$ excluding 13 to simulate a distribution shift. In the residual $\beta_{\ell}$ analysis, we use $K=13$ for in-distribution (ID) training and $K=17$ for out-of-distribution (OOD) training, allowing for a controlled comparison between generalization and memorization behavior. The model receives the first 9 tokens as input and is trained to predict the signal component of the 10th token.
@@
C.2 Finetuning Settings
We fine-tuned all layers end-to-end using AdamW (learning rate $5\times 10^{-6}$, weight decay $0.01$) with a linear schedule (warmup ratio $0.1$) and early stopping on validation loss. Training converged in all settings.
C.3 Residual Scaling with Learnable $\beta_{\ell}$ Parameters
@@
The in-distribution (ID) setting is using the training split of “gen_train23_test2to10”. Out-of-distribution (OOD) is conducted on held-out CLUTRR configurations that do not overlap with the training distribution: for GPT models we evaluate on the test split of “gen_train23_test2to10” (lengths 2 through 10), while for Qwen and LLaMA we use the test split of “rob_train_sup_23_test_all_23”, which defines a more challenging OOD setting by jointly varying both the clause length and the status of noise facts introduced into the story, given the higher modeling capacity of these models. In all cases, tasks 1.2 and 1.3 are excluded from the OOD sets, as they overlap with the training configuration and are therefore treated as in-distribution rather than out-of-distribution.
Training and Evaluation Settings
@@
GPT-2 Small 30.98 71.45
GPT-2 Medium 34.21 68.35
Qwen-2.5-0.5B 31.94 83.20 57.06
LLaMA-3.1-8B 57.68 94.00 77.39
We perform residual $\beta_{\ell}$ training on top of a model that has already been fine-tuned on the target dataset, keeping all previously learned weights frozen and optimizing only the $\beta_{\ell}$ parameters using HuggingFace’s Trainer API. We sweep over batch sizes from 4 to 32 depending on GPU memory constraints, and tune the learning rate between $5\times 10^{-5}$ and $5\times 10^{-6}$. We train for 3–10 epochs based on model capacity and task complexity. Optimization is performed using the AdamW optimizer with a weight decay of 0.01. Evaluation is conducted by comparing the next predicted token with the next ground-truth target token.
@@ Appendix E Full Quantitative Results with Confidence Estimates
We compute mutual information following the methodology in Appendix A. We subsample between 50 and 100 test examples per experiment and report results with statistical significance based on 5 seeds.
@@
Figure 9: Predictive information $I(Z;Y)$ across models and datasets. Lighter line colors represent later training epochs. Each curve shows the mean across 5 random seeds (0, 1, 2, 3, 42), and the shaded region denotes a 2-sigma (~96%) confidence interval.
@@
In addition to predictive information $I(Z^{(\ell)};Y)$, we compute the incremental information gain $I(\Delta Z^{(\ell)};Y)$ at each layer. This quantifies the information contribution for each transformer block.
@@
(a) GPT-2 Small on CLUTRR
(b) GPT-2 Small on Synthetic Arithmetic
(c) Qwen-2.5-0.5B on CLUTRR
(d) Qwen-2.5-0.5B on Synthetic Arithmetic
(e) LLaMA-3.1-8B on CLUTRR
(f) LLaMA-3.1-8B on Synthetic Arithmetic
Figure 12: Residual scaling coefficients $\beta_{\ell}$ across all transformer layers. Each curve shows the mean across five random seeds, and the shaded region denotes 1-sigma error bar. ID training emphasizes later layers, while OOD training has lower final layer weight. When the task is simple relative to the model’s capability, such as LLaMA on the synthetic task, the final-layer residual coefficients are similar for ID and OOD, consistent with Figure 9.

### v4

@@ 5.1 Predictive Information: Information Peaks at Intermediate Layers
Figure 2: Evolution of predictive information $I(Z_{\ell};Y)$, with lighter curves indicating later training steps. Each curve exhibits a three-phase trend: early layers rise, mid layers peak, and late layers decline.
@@
Information Peak (intermediate layers). $I(Z_{\ell};Y)$ continues to rise through the mid-to-upper layer, typically peaking before the final few blocks. For the GPT-2 Small on Synthetic Arithmetic dataset, the peak reaches $I\!\approx\!0.9$ in layers 10-11, coinciding with a jump from $\approx 0\%$ to $85\%$ ID accuracy and $59\%$ OOD accuracy. These layers appear to play a critical role in synthesizing abstract features that are essential for generalization.
@@
Layer $\mathbf{I(Z;Y)}$ Test Accuracy (%) All In-Dist. Out-Dist. Layer 1 0.0585 0.00 0.00 0.00 Layer 2 0.0673 0.00 0.00 0.00 Layer 3 0.0969 0.00 0.00 0.00 Layer 4 0.1880 0.00 0.00 0.00 Layer 5 0.3038 0.00 0.00 0.00 Layer 6 0.3991 0.00 0.00 0.00 Layer 7 0.5497 4.00 0.10 7.90 Layer 8 0.8067 19.30 0.00 38.60 Layer 9 0.8604 22.35 1.30 43.40 Layer 10 0.9044 57.00 54.30 59.70 Layer 11 0.8650 70.35 85.10 55.60 Layer 12 0.1783 71.95 93.80 50.10
@@
(a) 7 layers: Monotonic Increase
@@
Figure 3: Truncating GPT–2 to 7 layers removes the MI peak; 9–layer variants begin to exhibit a shallow peak, and the full 12–layer model shows a pronounced decline.
We empirically vary the depth of GPT-2 Small models fine-tuned on the CLUTRR dataset to examine how model capacity shapes information dynamics, shown in Figure 3. When the model is truncated to 7 layers, $I(Z;Y)$ increases monotonically—indicating insufficient capacity to develop an information peak. Upon increasing the depth to 9 layers, a shallow peak emerges, signifying the threshold at which the model begins to distinguish generalized features from memorized signals. The full 12-layer model exhibits a clear peak followed by a decline, confirming that the generalization ridge emerges only beyond a certain capacity threshold. These results underscore the role of architectural depth in information dynamics and sufficient capacity is required for the generalization ridge to emerge.
@@
This discussion aims to provide an intuitive illustration of our generalization ridge hypothesis, highlighting how attention to task-relevant signal tokens shifts across layers (the generalizable information). Specifically, we computed average signal attention across layers, identified the layer with peak signal attention, and compared it to final-layer signal and last-token attention, results are in Table 2. Signal tokens are defined as task-relevant tokens that the model must focus on to solve the task—for the synthetic dataset, these are tokens that appear after the character ‘S’; for CLUTRR, kinship-related entities; and for ECQA, the token corresponding to the correct answer option. This allows us to quantify where in the network attention to semantically important information is concentrated. Our findings show a pattern that supports our hypothesis: (1) Mid-to-late layers peak in signal attention, coinciding with the predictive information ridge (Figure 2), indicating where generalizable representations are strongest. (2) Final layers show reduced attention to signal tokens, suggesting a shift toward memorization to specific data point rather than predictive information abstraction. For example, in Qwen-2.5-0.5B on ECQA, signal attention peaks at Layer 17 (0.2307) but drops to 0.0104 in the final layer, where last-token attention dominates (0.5015).
@@
GPT-2 Small Synthetic 0.0222 0.0409 8 (0.0740) 0.0562 0.0348
GPT-2 Medium CLUTRR 0.0086 0.0086 18 (0.0270) 0.0124 0.1649
Qwen-2.5-0.5B ECQA 0.0225 0.0317 17 (0.2307) 0.0104 0.2169
LLaMA-3.1-8B ECQA 0.0220 0.0394 21 (0.1595) 0.0094 0.5015
@@ 5.2 Incremental Information Gain: Information Concentrates at Intermediate Layers
$z^{(\ell)}=z^{(\ell-1)}+\beta_{\ell}\cdot\text{block}^{(\ell)}(z^{(\ell-1)}),\qquad\beta_{\ell}\in\mathbb{R}.$
@@ E. Drive
Each synthetic sample consists of a sequence of 10 symbolic elements, where each element is formatted as S{signal}_N{noise}. The signal values follow an arithmetic progression modulo $K$, and the noise values are independently sampled from a uniform distribution with a fixed range of 100. During training, the modulus $K$ is set to 13. For evaluation, test sequences are generated using values of $K$ from the range $[5,26]$ excluding 13 to simulate a distribution shift. The model receives the first 9 tokens as input and is trained to predict the signal component of the 10th token.
@@
C.2 Residual Scaling with Learnable $\beta_{\ell}$ Parameters
@@
For Synthetic Arithmetic, the in-distribution (ID) setting is using $K=13$, the out-of-distribution (OOD) setting is using $K=[5,26]$ excluding 13. For CLUTRR, the in-distribution (ID) setting is using the “1.2” and “1.3” split of “gen_train23_test2to10”. Out-of-distribution (OOD) is conducted on held-out CLUTRR configurations that do not overlap with the training distribution: the “1.4” to “1.10” split of “gen_train23_test2to10”. For ECQA, we use training set as the ID setting and CLUTRR as the OOD setting.
C.3 Training and Evaluation Settings
@@
GPT-2 Small 56.57 71.95
GPT-2 Medium 62.11 73.15
Qwen-2.5-0.5B 58.61 88.80 65.45
LLaMA-3.1-8B 72.62 89.90 80.49
We finetune the model using HuggingFace’s Trainer API with batch sizes of 16, the learning rate of $\alpha=5\times 10^{-6}$, and AdamW optimizer with weight decay of 0.01. We perform residual $\beta_{\ell}$ training on top of a model that has already been fine-tuned on the target dataset, keeping all previously learned weights frozen and optimizing only the $\beta_{\ell}$ parameters. We use batch sizes of 4, the learning rate of 0.001, Adam optimizer and epochs of 10. Evaluation is conducted by comparing the next predicted token with the next ground-truth target token.
@@ Appendix E Full Quantitative Results with Confidence Estimates
We subsample 100 test examples per experiment to compute mutual information and report results with statistical significance based on 5 seeds.
@@
Figure 9: Predictive information $I(Z;Y)$ across models and datasets. Lighter line colors represent later training steps. Each curve shows the mean across 5 random seeds (0, 1, 2, 3, 42), and the shaded region denotes a 2-sigma (~96%) confidence interval.
@@
In addition to predictive information $I(Z^{(\ell)};Y)$, we compute the incremental information gain $I(\Delta Z^{(\ell)};Y)$ for each transformer block.
@@
(a) GPT-2 Small on Synthetic Arithmetic
(b) GPT-2 Medium on CLUTRR
(c) Qwen-2.5-0.5B on ECQA
(d) LLaMA-3.1-8B on ECQA
Figure 12: Residual scaling coefficients $\beta_{\ell}$ across all transformer layers. Each curve shows the mean across five random seeds, and the shaded region denotes 1-sigma error bar. ID training emphasizes later layers, while OOD training has lower final layer weight.